# Filtered Direct Preference Optimization

### v3

@@ 2 Related work
We examine methods for aligning LMs with human preferences, focusing on RLHF and its alternatives. Most RLHF approaches utilize an RM (Ouyang et al., 2022; Touvron et al., 2023; Dubois et al., 2023; Casper et al., 2023). These methods fine-tune LMs using RL algorithms such as REINFORCE (Williams, 1992; Rennie et al., 2017), proximal policy optimization (PPO) (Schulman et al., 2017), or their variants Sutton and Barto (2018). However, there are notable reinforcement-learning-free approaches (Zhao et al., 2023; Liu et al., 2024), and learning-free methods that leverage the RM at decoding time, with best-of-N (BoN) sampling being a prominent example (Stiennon et al., 2020; Nakano et al., 2021).
@@
DPO and its extensions (Azar et al., 2023; Tang et al., 2024; Pal et al., 2024; Singh et al., 2024) represent significant RM-free methods. Some DPO variants explore different regularizations (Wang et al., 2024) or use a divided dataset for stepwise training (Gou and Nguyen, 2024). Other variants propose adapting DPO online (Xu et al., 2023; Guo et al., 2024) or evaluating the quality difference between chosen and rejected responses for adding an offset to the DPO objective function (Amini et al., 2024) or incorporating curriculum learning (Gou and Nguyen, 2024). These approaches focus on response quality, which is relevant to our method.
@@ 4 Filtered Direct Preference Optimization
5:Initialize filtered-preference dataset $\mathcal{D}_{\rm f}:=\mathcal{D}_{\rm pref}$, epoch number $m:=0$.
@@
10:              Discard $(x,y_{c},y_{r})$ from $\mathcal{D}_{\rm f}$.
@@
13:     Update LM $\pi_{\theta}$ on $\mathcal{D}_{\rm f}$ for one epoch using DPO.
14:     Increment epoch number $m:=m+1$.
15:end while
16:return Optimized LM $\pi_{\theta}$.
@@ 4.2 Background and Motivation for fDPO
$=\sigma\!\bigg{(}\!\beta\log\frac{\pi_{\theta}(y_{c}\,|\,x)}{\pi_{\rm ref}(y_{c}\,|\,x)}-\beta\log\frac{\pi_{\theta}(y_{r}\,|\,x)}{\pi_{\rm ref}(y_{r}\,|\,x)}\!\bigg{)}.$
@@ 5.2 Datasets
We provide the evaluation scores of the gold RM for these datasets in Table 4 in Appendix C.
@@ 5.3 Effect of Data Quality to Performance of RLHF and DPO
Figure 1 (A) shows the results, where the mean and standard error of the gold reward with five independent runs are presented. Notably, while DPO experienced a decline in efficacy when trained on the mixed-quality dataset relative to the high-quality one, RLHF showed an intriguing resilience, sustaining comparable performance levels across both datasets. This differential impact starkly highlights the greater susceptibility of DPO to dataset quality, suggesting that the RM-based approach, including fDPO, may offer more stable performance when the preference dataset quality cannot be consistently assured. However, RLHF’s overall gold reward was lower than DPO’s. Therefore, subsequent experiments focus on DPO.
@@ 5.4 Evaluation of fDPO on AlpacaFarm dataset
Figure 2 present the results of DPO and fDPO. The results shows that the performance of DPO trained on the high-quality dataset and fDPO trained on the mixed-quality dataset were on par. It indicates that fDPO has successfully circumvented the performance decline typically observed with DPO, thereby showcasing its potential to improve DPO performance where dataset quality is inconsistent. Corresponding learning curves are included in Appendix C.
@@ 5.5 Evaluation of fDPO under Realistic RLHF Settings on Anthropic HH Datasets
The evaluation metrics used in this study included the gold RM score, as described in the previous sections, and an additional evaluation using GPT-4o to determine the win rate. The win rate indicates how often responses generated by the trained LM were preferred over those generated by the initial SFT model.
As shown in Table 1, based on three independent runs, fDPO outperformed the baseline in both evaluation metrics of the gold RM scores and GPT-4o win rates. The superior GPT-4o evaluation results suggest that fDPO is not merely optimizing for the reward model but is also learning to generate higher-quality responses from a human evaluation perspective. This demonstrates the effectiveness of our approach under realistic RLHF settings, providing a viable solution for scenarios where high-quality responses are limited.
@@ 7 Limitations
The fDPO method shows promise, but it has some limitations. First, the method requires a reward model, which might be a drawback as it increases the complexity and computational time of the method. However, the availability of high-quality reward models provides an opportunity to leverage these high-end models within the DPO framework. Exploring the use of implicit rewards in DPO instead of an explicit reward model could also address some complications associated with training a separate reward model. Second, the algorithm is implemented in its simplest form, suggesting significant room for improvement and optimization. Third, our approach does not account for rejected responses, which could further enhance performance if considered. Finally, our experiments are limited to relatively small LLMs and comparisons with DPO. Future work should explore combining fDPO with other DPO-related extensions and conducting comparisons with other RLHF methods, especially with larger LLMs.
@@ Appendix B Justification on filtering chosen responses
  • the magnitudes of the gradients for $\log\pi_{\theta}(y_{c})$ and $\log\pi_{\theta}(y_{r})$ are similar, i.e.,
@@
  • the gradients for $\log\pi_{\theta}(y_{c})$ and $\log\pi_{\theta}(y_{r})$ are nearly orthogonal, i.e.,
    $\nabla_{\theta}\log\pi_{\theta}(y_{c})^{\!\top}\nabla_{\theta}\log\pi_{\theta}(y_{r})\simeq 0,$
@@
where $\alpha$ is the learning rate and is sufficiently small, the sensitivity of $\pi_{\theta}(y_{c})$, defined as the magnitude of change in probability, $\Delta\pi_{\theta}(y)$, is approximately $\delta$ times higher than that of $\pi_{\theta}(y_{r})$.
@@
Given the assumptions, the magnitudes of the gradients for $\log\pi_{\theta}(y_{c})$ and these gradients are nearly orthogonal. Hence, the impact of $\Delta\theta$ on $\log\pi_{\theta}(y_{c})$ and $\log\pi_{\theta}(y_{r})$ would be similar in magnitude but differ in direction. However, due to the ratio $\pi_{\theta}(y_{c})/\pi_{\theta}(y_{r})=\delta$, the rate of change in $\pi_{\theta}(y_{c})$ is amplified by a factor of $\delta$ compared to $\pi_{\theta}(y_{r})$. Thus, under the DPO update, $\pi_{\theta}(y_{c})$ demonstrates a sensitivity that is approximately $\delta$ times higher than that of $\pi_{\theta}(y_{r})$. ∎
@@ C.1 Hyperparameters
epochs 8 (16 for fDPO) 4 (8 for fDPO) 3 4 (8 for fDPO)
batch size 64 64 32 64
learning_rate $1\times 10^{-6}$ $2\times 10^{-7}$ $1\times 10^{-6}$ $5\times 10^{-6}$
@@ C.3 Additional results on AlpacaFarm dataset
C.3.1 Detailed evaluation
The evaluation scores of the gold reward model for the preference datasets (high-quality, low-quality, mix-quality) of the AlpacaFarm dataset are detailed in Table 4. The mix-quality datasets (Mix 1-5) each consist of 50% randomly sampled data from the high-quality dataset and the low-quality dataset, using random seeds 1-5, respectively.
@@ C.3.1 Detailed evaluation
Figure 6 provides the learning curves for DPO and fDPO with the the 160M-sized LM, corresponding to the final performances depicted in Figure 1 (B) of the main text. The curves show that even though fDPO processes double the number of epochs compared to DPO, the total number of steps for fDPO is fewer than that for DPO. This reduction is due to the filtering process of fDPO, which decreased the data over epochs, resulting in fewer steps per epoch, as demonstrated in Figure 8. Additionally, when assessed using KL divergence, the performance of fDPO shows a trend towards converging with the DPO trained on the high-quality dataset, suggesting that fDPO can reduce the performance gap even when trained on mixed-quality data.
Figure 7 presents the learning curves for DPO and fDPO applied to the mix-quality dataset with the 1.4B LM and the low-quality dataset with the 160M LM. In both contexts, fDPO consistently improved the performance of DPO over steps, echoing the results observed in the mix-quality dataset scenario with the 160M LM.
@@
C.3.2 Analysis of filtered samples of fDPO
@@ C.4 Additional results on Anthropic HH Datasets
C.4.1 Detailed evaluation
The evaluation scores of the gold reward model for our preference datasets (original, SFT-model-generated, mix-quality) of the Anthropic HH dataset are detailed in Table 5. The mix-quality datasets (Mix 1-3) each consist of 25% randomly sampled responses from the original Anthropic HH dataset and 75% from the responses generated by the SFT model, using random seeds 1-3, respectively.
@@ C.4.1 Detailed evaluation
Figure 9 provides using the helpful dataset and the harmless dataset of the Anthropic HH datasets, respectively. The curves show that even though fDPO processes double the number of epochs compared to DPO, the total number of steps for fDPO is fewer due to the filtering process.
@@
C.4.2 Future Evaluation
In addition to the aforementioned experiment, we conduct an experiment with 50% original and 50% SFT-generated responses. The detailed results of this experiment are provided in Table 6.

### v4

@@ 2 Related work
We examine methods for aligning LMs with human preferences, focusing on RLHF and its alternatives. Most RLHF approaches utilize an RM (Ouyang et al., 2022; Touvron et al., 2023; Dubois et al., 2023; Casper et al., 2023). These methods fine-tune LMs using RL algorithms such as REINFORCE (Williams, 1992; Rennie et al., 2017), proximal policy optimization (PPO) (Schulman et al., 2017), or their variants Sutton and Barto (2018). However, there are notable reinforcement-learning-free approaches (Zhao et al., 2023; Liu et al., 2024a), and learning-free methods that leverage the RM at decoding time, with best-of-N (BoN) sampling being a prominent example (Stiennon et al., 2020; Nakano et al., 2021).
@@
DPO and its extensions (Azar et al., 2023; Tang et al., 2024; Pal et al., 2024; Singh et al., 2024) represent significant RM-free methods. Some DPO variants explore different regularizations (Wang et al., 2024) or use a divided dataset for stepwise training (Gou and Nguyen, 2024). Other variants propose adapting DPO online (Xu et al., 2023; Guo et al., 2024), where an RM is used to evaluate newly generated training data. While both online DPO and fDPO employ RMs, fDPO remains an offline approach that filters the dataset and does not use generated data for training. This distinction makes fDPO less reliant on the accuracy of RMs, as it focuses solely on data refinement rather than reward optimization. Liu et al. (2024b), on the other hand, utilizes an RM for rejection sampling to adjust the distribution of the preference data by aligning it with the optimal LM distribution. In contrast, fDPO focuses on filtering low-quality data. Other DPO variants evaluate the quality difference between chosen and rejected responses for adding an offset to the DPO objective function (Amini et al., 2024) or incorporating curriculum learning (Gou and Nguyen, 2024). These approaches focus on response quality, which is relevant to our method.
@@ 4 Filtered Direct Preference Optimization
[…]
@@
10:              Discard $(x,y_{c},y_{r})$ from $\mathcal{D}_{\rm filtered}$.
@@
13:     Update preference data $\mathcal{D}_{\rm pref}:=\mathcal{D}_{\rm filtered}$
14:     Update LM $\pi_{\theta}$ on $\mathcal{D}_{\rm pref}$ for one epoch using DPO.
15:     Increment epoch number $m:=m+1$.
16:end while
17:return Optimized LM $\pi_{\theta}$.
@@ 4.2 Background and Motivation for fDPO
$=\sigma\!\bigg{(}\!\beta\log\frac{\pi_{\theta}(y_{r}\,|\,x)}{\pi_{\rm ref}(y_{r}\,|\,x)}-\beta\log\frac{\pi_{\theta}(y_{c}\,|\,x)}{\pi_{\rm ref}(y_{c}\,|\,x)}\!\bigg{)}.$
@@ 5.2 Datasets
We provide the evaluation scores of the gold RM for these datasets in Table 4 (Appendix C.3.1).
@@ 5.3 Effect of Data Quality to Performance of RLHF and DPO
Figure 1 (A) shows the results, where the mean and standard error of the gold reward with five independent runs are presented. Notably, while DPO experienced a decline in efficacy when trained on the mixed-quality dataset relative to the high-quality one, RLHF showed an intriguing resilience, sustaining comparable performance levels across both datasets. This differential impact starkly highlights the greater susceptibility of DPO to dataset quality, suggesting that the RM-based approach, including fDPO, may offer more stable performance when the preference dataset quality cannot be consistently assured. However, RLHF’s overall gold reward was comparable to DPO’s. Nevertheless, our focus remains on offline alignment, as offline alignment methods could be complementary to online alignment approaches (e.g., applying an offline method first, followed by an online one). Therefore, subsequent experiments focus on DPO.
@@ 5.4 Evaluation of fDPO on AlpacaFarm dataset
Figure 2 present the results of DPO and fDPO. The results shows that the performance of DPO trained on the high-quality dataset and fDPO trained on the mixed-quality dataset were on par. It indicates that fDPO has successfully circumvented the performance decline typically observed with DPO, thereby showcasing its potential to improve DPO performance where dataset quality is inconsistent. Corresponding learning curves are included in Appendix C.3.2.
@@ 5.5 Evaluation of fDPO under Realistic RLHF Settings on Anthropic HH Datasets
The evaluation metrics used in this study included the gold RM score, as described in the previous sections, and an additional evaluation using GPT-4o to determine the win rate. The win rate indicates how often responses generated by the trained LM were preferred over those generated by the initial SFT model. Additionally, Appendices C.4.4 and C.4.6 provide small-scale human evaluations and qualitative analyses of the generated responses.
Table 1 shows the results of the evaluations based on three independent runs. fDPO outperformed the baseline in both evaluation metrics: the gold RM scores and GPT-4o win rates. The superior GPT-4o evaluation results suggest that fDPO is not merely optimizing for the reward model but is also learning to generate higher-quality responses from a human evaluation perspective. This demonstrates the effectiveness of our approach under realistic RLHF settings, providing a viable solution for scenarios where high-quality responses are limited.
@@ 7 Limitations
The fDPO method shows promise, but it has some limitations. First, the method requires a reward model, which might be a drawback as it increases the complexity and computational time of the method. However, the availability of high-quality reward models provides an opportunity to leverage these high-end models within the DPO framework, though in more specialized tasks such as legal arguments or mathematical problems, developing task-specific reward models may be necessary. Exploring the use of implicit rewards in DPO instead of an explicit reward model could also address some complications associated with training a separate reward model. Second, the algorithm is implemented in its simplest form, suggesting significant room for improvement and optimization. For example, future work could explore alternative strategies such as data replacement or incorporating curriculum learning based on sample quality. Furthermore, our approach does not account for rejected responses, which could further enhance performance if considered. Finally, our experiments are limited to relatively small LLMs and comparisons with DPO. Future work should explore combining fDPO with other DPO-related extensions and conducting comparisons with other RLHF methods, especially with larger LLMs.
@@ Appendix B Justification on filtering chosen responses
  • the L2-norms of the gradients for $\log\pi_{\theta}(y_{c})$ and $\log\pi_{\theta}(y_{r})$ are similar, i.e.,
@@
  • the normalized gradients for $\log\pi_{\theta}(y_{c})$ and $\log\pi_{\theta}(y_{r})$ are nearly orthogonal, i.e.,
    $\frac{\nabla_{\theta}\log\pi_{\theta}(y_{c})^{\!\top}\nabla_{\theta}\log\pi_{\theta}(y_{r})}{\|\nabla_{\theta}\log\pi_{\theta}(y_{c})\|\|\nabla_{\theta}\log\pi_{\theta}(y_{r})\|}\simeq 0,$
@@
where $\alpha$ is the learning rate and is sufficiently small, the sensitivity of $\pi_{\theta}(y_{c})$, defined as the magnitude of change in probability, is approximately $\delta$ times higher than that of $\pi_{\theta}(y_{r})$.
@@
Given the assumptions, the norms of the gradients for $\log\pi_{\theta}(y_{c})$ and $\log\pi_{\theta}(y_{r})$ are similar, and these normalized gradients are nearly orthogonal. Hence, the impact of $\Delta\theta$ on $\log\pi_{\theta}(y_{c})$ and $\log\pi_{\theta}(y_{r})$ would be similar in magnitude but differ in direction. However, due to the ratio $\pi_{\theta}(y_{c})/\pi_{\theta}(y_{r})=\delta$, the rate of change in $\pi_{\theta}(y_{c})$ is amplified by a factor of $\delta$ compared to $\pi_{\theta}(y_{r})$. Thus, under the DPO update, $\pi_{\theta}(y_{c})$ demonstrates a sensitivity that is approximately $\delta$ times higher than that of $\pi_{\theta}(y_{r})$. ∎
@@
Experimental validation of assumptions:
[…]
Figure 6 shows the results of this experiment. The results suggest that the assumptions in Proposition B.1 largely hold, with the log norm ratio of the gradients and the cosine similarity between them being close to the expected value of zero. Furthermore, the log norm ratio and cosine similarity show minimal dependence on $\delta=\pi(y_{i}|x)/\pi(y_{j}|x)$.
Figure 6: Results of the empirical validation of the assumptions in Proposition B.1. The plots display the log delta $\log(\pi(y_{i}|x)/\pi(y_{j}|x))$, log norm ratio $\log(\|\nabla\log\pi(y_{i}|x)\|/\|\nabla\log\pi(y_{j}|x)\|)$, and cosine similarity $\cos(\nabla\log\pi(y_{i}|x),\nabla\log\pi(y_{j}|x))$ for 120 pairs of responses generated by GPT-2 large. The results show that the assumptions largely hold, with minimal dependence on $\delta=\pi(y_{i}|x)/\pi(y_{j}|x)$.
@@ C.1 Hyperparameters
epochs 8 (16 for fDPO) 4 (8 for fDPO) 8 4 (8 for fDPO)
batch size 64 64 128 64
learning_rate $1\times 10^{-6}$ $2\times 10^{-7}$ $4\times 10^{-6}$ $5\times 10^{-6}$
@@ C.3 Additional results on AlpacaFarm dataset
C.3.1 Dataset evaluation
The evaluation scores of the gold reward model for the preference datasets (high-quality, low-quality, mix-quality) of the AlpacaFarm dataset are detailed in Table 4. The mix-quality datasets (Mix 1-5) each consist of 50% randomly sampled data from the high-quality dataset and the low-quality dataset, using random seeds 1-5, respectively. As expected, the evaluation scores follow a clear trend where high-quality datasets achieve the highest scores, followed by the mix-quality datasets, and finally the low-quality datasets. This trend is consistent across both model sizes (160M and 1.4B).
@@ C.3.1 Dataset evaluation
C.3.2 Learning curves
Figure 7 provides the learning curves for DPO and fDPO with the 160M-sized LM, corresponding to the final performances depicted in Figure 1 (B) of the main text. The curves indicate that although fDPO processes twice the number of epochs compared to DPO, the total number of steps for fDPO is fewer due to its filtering process. As data is filtered over epochs, the number of steps per epoch decreases, as demonstrated in Figure 9 (unfiltered ratio). Moreover, when evaluated based on KL divergence, fDPO’s performance converges towards that of DPO trained on the high-quality dataset, suggesting that fDPO can effectively close the performance gap even when trained on mixed-quality data.
Figure 8 illustrates the learning curves for DPO and fDPO using the mix-quality dataset with the 1.4B LM and the low-quality dataset with the 160M LM. In both cases, fDPO consistently outperforms DPO as training progresses, mirroring the trends observed in the mix-quality dataset scenario with the 160M LM.
@@ C.3.2 Learning curves
C.3.3 Analysis of filtered samples of fDPO
@@ C.4 Additional results on Anthropic HH Datasets
C.4.1 Dataset Evaluation
The evaluation scores of the gold reward model for our preference datasets (original, SFT-model-generated, and mix-quality) of the Anthropic HH dataset are shown in Table 5. The mix-quality datasets (Mix 1-3) each consist of 25% randomly sampled responses from the original Anthropic HH dataset and 75% from the responses generated by the SFT model, using random seeds 1-3, respectively.
The table shows that both the helpful and harmless datasets follow a similar trend: the original dataset achieves the highest scores, followed by the mix-quality dataset, with the SFT-model-generated dataset scoring the lowest. The mix-quality datasets exhibit intermediate scores, which is expected given that they are a mixture of the original and SFT-model-generated datasets. Notably, the higher scores of the original dataset compared to the SFT-model-generated dataset suggest that the original data is of higher quality.
@@ C.4.1 Dataset Evaluation
C.4.2 Learning curves
Figure 10 shows the learning curves for DPO and fDPO over steps using the Anthropic HH datasets. The results indicate that although fDPO is set to process double the number of epochs compared to DPO, the total number of steps for fDPO is fewer due to the filtering process, which reduces the number of samples per epoch over time.
@@ C.4.2 Learning curves
C.4.3 Experiment with different data mixes
In addition to the previous experiment, we conducted another experiment using a data mix consisting of 50% original responses and 50% SFT-generated responses. The detailed results of this experiment are shown in Table 6. The table shows that fDPO outperforms DPO in both the helpful and harmless datasets. This result is consistent with the findings from Table 1, which used a mix of 75% original and 25% SFT-generated responses.